OpenStudy (anonymous):

In circle A, AED = 48 degrees and DAE = 44 degrees. Determine if AC = EF. Explain why or why not and show all work.

OpenStudy (anonymous):

Here is the circle for the question

OpenStudy (anonymous):

@mww @welshfella care to help on this one

OpenStudy (mww):

This one is simple. But the question is not to show AC = EF, rather that they are perpendicular to each other. To do this, show that the angle ADE is a right angle (90 degrees) So look at triangle ADE. What are the size of the angles inside ADE you have been given and what must all angles inside a triangle add up to? This will help you figure out angle ADE. It has to be 90 for AC to be perpendicular to EF.

OpenStudy (anonymous):

@mww they all have to equal to 180 am i correct?

OpenStudy (phi):

the 3 angles in a triangle add up to 180 if you are given 2 of the angles, you can figure out the 3rd. Do you see how ?

OpenStudy (anonymous):

sort of, ive never been good at any type of math

OpenStudy (phi):

you can try on this. what do the 2 angles they give you add up to ?

OpenStudy (anonymous):

they equal up to 92 degrees

OpenStudy (phi):

yes. if we call the 3rd angle x (x for unknown) 92 + x = 180 which is a way to write: 92 plus an unknown amount add up to 180

OpenStudy (phi):

maybe it's not clear how to find that number? if you had 1+ ? = 2 you know what number that is , right ?

OpenStudy (anonymous):

it would be 1, because everyone know 1 + 1 = 2

OpenStudy (phi):

yes. and 1+ ? = 3 the ? would be 2. it turns out if we do 3-1 , we get the answer in your problem 92+x = 180 x = 180 - 92

OpenStudy (anonymous):

okay i totally made this harder on myself when its actually pretty easy wow

OpenStudy (phi):

what do you get for the 3rd angle ?

OpenStudy (anonymous):

i got 88 degrees

OpenStudy (phi):

yes. next, what angle do two perpendicular lines make ? any idea ?

OpenStudy (anonymous):

the form a 90 degrees angle

OpenStudy (phi):

yes, and 88 is not 90 so the two lines AC and EF are *not* perpendicular

OpenStudy (anonymous):

i feel s stupid hah that was such an easy question but i just made it harder on myself. Thank you so much!

OpenStudy (phi):

If this stuff was obvious, you would not have to be taught. One reason people are taught math is to practice "thinking" which is a hard thing to learn. (but useful!)